# URS: A Unified Neural Routing Solver for Cross-Problem Zero-Shot Generalization

### v1

@@ A.2 Neural Routing Solvers with Multi-Task Learning
To address the challenge of cross-problem generalization, growing attention has shifted to multi-task learning capable of adapting to diverse routing problems. Existing efforts can broadly fall into two categories: (1) constraint combination-based multi-task learning and (2) adapter-based fine-tuning frameworks. The former line (e.g., MTPOMO (Liu et al., 2024a)) treats VRP variants as combinations of a predefined set of attributes and trains a shared backbone, achieving promising results on up to $16$ VRP variants. Follow-up work extends this paradigm by introducing mixture-of-experts (MoE) modules (Zhou et al., 2024b), heavy decoder designs plus knowledge distillation (Zheng et al., 2025a), and constraint-aware dual-attention (Li et al., 2025a) or updated Transformer structure (Berto et al., 2024) to improve model performance. While this enables knowledge sharing across problems, its coverage is inherently bounded by the hand-specified attribute set. The second family trains a shared backbone and performs adaptation to each variant via lightweight adapters. For instance, Lin et al. (2024) pretrain a TSP backbone and adapt to new variants through adapter-based fine-tuning. Drakulic et al. (2025) integrate edge features and adopt supervised multi-task pretraining for obtaining a powerful backbone before fine-tuning on novel tasks. Wang et al. (2025a) propose a multi-armed bandit strategy to realize more efficient multi-task training. Although these reduce retraining cost, they fail to achieve zero-shot generalization. In contrast, we focus on a constructive neural routing solver supported with zero-shot capability: once trained, it can directly generate high-quality solutions for previously unseen problems without any fine-tuning.
A.3 LLM-based Algorithm Generation in VRPs
Employing LLMs to solve VRPs is a new research direction. As an early attempt, Yang et al. (2023) utilizes the LLM as a direct optimizer, prompting an LLM to directly generate solutions for TSP instances. A more predominant stream leverages LLMs as automated algorithm designers to generate algorithms for solving VRPs. These methods often integrate LLMs with a search procedure. EoH (Liu et al., 2024b) combined an LLM with Evolutionary Computation to design heuristic components for a Guided Local Search (GLS) framework. ReEvo (Ye et al., 2024a) incorporated a reflection mechanism to refine heuristics for iterative solution construction. HSEvo (Dat et al., 2025) additionally introduced a diversity mechanism for designing heuristics solving TSP and OP. And AHD-MCTS (Zheng et al., 2025b) employed Monte Carlo Tree Search (MCTS) to guide the design of heuristics for solving TSP and CVRP. Despite their promise, these methods share common limitations rooted in their reliance on the LLM’s implicit, pre-trained knowledge as the core design logic. The process of discovering effective algorithms demands extensive search procedures and is computationally expensive. Furthermore, the heuristics generated by the LLM are susceptible to overfitting to the specific instances used during the design phase, limiting their generalization capabilities. In contrast, our URS overcomes these challenges by establishing a clear division of labor: the LLM is tasked with satisfying complex constraints, while our URS model is responsible for ensuring solution quality. This approach significantly reduces the reasoning burden on the LLM, mitigating the risk of overfitting and enabling URS to effectively solve a more extensive range of VRPs.
@@ Appendix B Setups of VRP Variants
In this paper, we consider over 100 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD). The node coordinates are randomly sampled from the unit square when available (in symmetric geometric cases, i.e., $d_{ij}=d_{ji}$), following AM (Kool et al., 2019). Below, we provide a comprehensive description of the data generation process for each constraint.
@@
The standard CVRP involves vehicles delivering goods from a depot to a set of customers with positive demands. The introduction of backhauls extends this formulation to the CVRP with Backhauls (CVRPB), where vehicles must also load goods from backhaul customers and return them to the depot, in addition to serving the standard linehaul (delivery) customers. Following the experimental setup of MTPOMO (Liu et al., 2024a), we generate customer demands by sampling from a discrete uniform distribution over the integers $\{1,\ldots,9\}$. Subsequently, 20% of these nodes are randomly designated as backhaul customers, and their corresponding demand values are negative, while the remainder are classified as linehaul customers. Notably, no precedence constraints are imposed between the servicing of linehaul and backhaul customers.To ensure the feasibility of the solutions constructed by our neural solver, we adhere to the MVMoE (Zhou et al., 2024b) framework’s policy that the first customer visited must be a linehaul node.
@@
Following MTPOMO (Liu et al., 2024a), we impose a duration limit (i.e., the maximum length of each vehicle route) of 3 in symmetric cases. Since the node coordinates are randomly sampled from the unit square, the maximum possible distance between any two nodes is $\sqrt{2}$. The setting is sufficient to ensure a vehicle can deliver goods to at least one customer and return to the depot, thus guaranteeing the feasibility of the solution. For the setting of asymmetric cases, we provide details in the description of the following asymmetric instances.
@@
Consistent with the MVMoE (Zhou et al., 2024b), we set time windows of the depot to $e_{0},l_{0}]$=[0,3] and its service time to 0. The service time $s_{i}$ for each customer node is set to 0.2. For the time windows for each customer node, we generate them following the methodology outlined in MVMoE (Zhou et al., 2024b).
@@
The generation of coordinates for the multiple depots aligns with that of other nodes, with the number of depots set to 3 (Berto et al., 2024). A vehicle is required to return to the same depot it departed from. To guarantee that each starting node is considered in this multi-depot variant, we set the number of trajectories for an instance to be the product of the number of depots and the number of customer nodes, maximizing the exploration of high-quality solutions.
@@
Under the Prize Collecting condition, each node (excluding the depot) is assigned both a prize and a penalty. A prize is collected for each visited node, while a penalty is incurred for each unvisited node. The objective is to minimize the total travel distance plus the sum of penalties for all unvisited nodes. A key constraint is that a minimum amount of total prize must be collected before the vehicle is permitted to return to the depot and terminate its route. Consistent with the AM (Kool et al., 2019) framework, the prizes and penalties are generated as follows: $prize\sim\text{Uniform}\left(0,\frac{4}{n}\right)$, and $penalty\sim\text{Uniform}\left(0,3\cdot\frac{k_{n}}{n}\right)$, where $n=100$ and $k_{n}=4$ in our experiments.
@@ Appendix C Unified Data Representation
In this section, we provide a detailed description of our UDR $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$ (see Table 5). Each node $\mathbf{u}_{i}$ consists of three components, i.e., $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$. where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a position identifier, unified attribute set, and node-type indicator, respectively. To distinguish linehaul/delivery from backhaul/pickup nodes, we assign negative demand values to backhaul/pickup nodes (positive for linehaul/delivery), making demand the only signed component, following  (Liu et al., 2024a; Zhou et al., 2024b).
@@ Appendix D Multi-hot Problem Representation
As shown in Table 6, we train URS on eleven classic and varied routing problems to ensure that each feature in the UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$. Then, we evaluate generalization on unseen problems on different combinations of ten constraint attributes (see Appendix B), and the detailed representations of each unseen problem can be found in Table 7 (documented VRP variants) and Table 8 (undocumented VRP variants).
@@
ACVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Table 7: The problem representations of $44$ VRP variants that have been extensively explored in current literature (Liu et al., 2024a; Berto et al., 2024). Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
@@
Table 8: The problem representations of $52$ VRP variants that have not yet been explored in current literature. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
@@
Appendix E Model Architecture
@@ E.1 Encoder
Let ${H}^{(\ell-1)}=\{\mathbf{h}_{i}^{(\ell-1)}\}_{i=0}^{n}$ denote the input to the $\ell$-th attention layer for $\ell=1,\ldots,L$. The outputs for the $i$-th node are computed as follows:
@@
where $\mathrm{IN}(\cdot)$ denotes instance normalization (Ulyanov et al., 2016), $\mathrm{MBM}$ represents the adopted mixed bias module (see Equation 5 for details), and $\mathrm{FF}(\cdot)$ in Equation 11 corresponds to a fully connected neural network with ReLU activation.
After $L$ attention layers, the final node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$ encapsulate the advanced feature representations of each node.
Notably, to ensure dimensional consistency across bias channels, we normalize $\bm{D}$ and $\bm{D^{\mathrm{T}}}$ by dividing their maximum value before passing the MBM.
@@ E.2 Decoder
$\bm{H}_{\text{dec}}(\bm{\lambda})=((\bm{\lambda}W_{1}+\mathbf{b}_{1})W_{2}+\mathbf{b}_{2})W_{3}+b_{3},$ (12)
where $W_{1}\in\mathbb{R}^{|\lambda|\times d_{h}},\ W_{2}\in\mathbb{R}^{d_{h}\times d_{h}},\ W_{3}\in\mathbb{R}^{d_{h}\times(5\times|\bm{\lambda}|)},\ \mathbf{b}_{1}\in\mathbb{R}^{d_{h}},\ \mathbf{b}_{2}\in\mathbb{R}^{d_{h}},\mathbf{b}_{2}\in\mathbb{R}^{(5\times|\bm{\lambda}|)}$ are learnable parameters, $d_{h}$ is $256$ in this work, following Lin et al. (2022). Next, we use five different linear projections to map the new representation $\bm{h}_{\text{dec}}^{i}(\bm{\lambda})$ of each row to the corresponding decoder parameters. The computation can be expressed as:
@@
where $W^{1}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{2}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{3}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(1\times d)}$, $W^{4}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{5}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$ are learnable parameters. In this way, we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate used linear projection matrices $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ conditioned on the multi-hot problem representation $\bm{\lambda}$, enabling the adaptive construction of high-quality solutions for different problems.

### v2

@@ A.2 Neural Routing Solvers with Multi-Task Learning
To address the challenge of cross-problem generalization, growing attention has shifted to multi-task learning capable of adapting to diverse routing problems. Existing efforts can broadly fall into two categories: (1) constraint combination-based multi-task learning and (2) adapter-based fine-tuning frameworks. The former line (e.g., MTPOMO (Liu et al., 2024)) treats VRP variants as combinations of a predefined set of attributes and trains a shared backbone, achieving promising results on up to $16$ VRP variants. Follow-up work extends this paradigm by introducing mixture-of-experts (MoE) modules (Zhou et al., 2024), heavy decoder designs plus knowledge distillation (Zheng et al., 2025), and constraint-aware dual-attention (Li et al., 2025a) or updated Transformer structure (Berto et al., 2025) to improve model performance. While this enables knowledge sharing across problems, its coverage is inherently bounded by the hand-specified attribute set. The second family trains a shared backbone and performs adaptation to each variant via lightweight adapters. For instance, Lin et al. (2024) pretrain a TSP backbone and adapt to new variants through adapter-based fine-tuning. Drakulic et al. (2025) integrate edge features and adopt supervised multi-task pretraining for obtaining a powerful backbone before fine-tuning on novel tasks. Wang et al. (2025a) propose a multi-armed bandit strategy to realize more efficient multi-task training. Although these reduce retraining cost, they fail to achieve zero-shot generalization. In contrast, we focus on a constructive neural routing solver supported with zero-shot capability: once trained, it can directly generate high-quality solutions for previously unseen problems without any fine-tuning.
@@ Appendix B Setups of VRP Variants
In this paper, we consider 110 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD). The node coordinates are randomly sampled from the unit square when available (in symmetric geometric cases, i.e., $d_{ij}=d_{ji}$), following AM (Kool et al., 2019). Below, we provide a comprehensive description of the data generation process for each constraint.
@@
The standard CVRP involves vehicles delivering goods from a depot to a set of customers with positive demands. The introduction of backhauls extends this formulation to the CVRP with Backhauls (CVRPB), where vehicles must also load goods from backhaul customers and return them to the depot, in addition to serving the standard linehaul (delivery) customers. Following the experimental setup of MTPOMO (Liu et al., 2024), we generate customer demands by sampling from a discrete uniform distribution over the integers $\{1,\ldots,9\}$. Subsequently, 20% of these nodes are randomly designated as backhaul customers, and their corresponding demand values are negative, while the remainder are classified as linehaul customers. Notably, no precedence constraints are imposed between the servicing of linehaul and backhaul customers. To ensure the feasibility of the solutions constructed by our neural solver, we adhere to the MVMoE (Zhou et al., 2024) framework’s policy that the first customer visited must be a linehaul node.
@@
Following MTPOMO (Liu et al., 2024), we impose a duration limit (i.e., the maximum length of each vehicle route) of 3 in symmetric cases. Since the node coordinates are randomly sampled from the unit square, the maximum possible distance between any two nodes is $\sqrt{2}$. The setting is sufficient to ensure a vehicle can deliver goods to at least one customer and return to the depot, thus guaranteeing the feasibility of the solution. For the setting of asymmetric cases, we provide details in the description of the following asymmetric instances.
@@
Consistent with the MVMoE (Zhou et al., 2024), we set time windows of the depot to $e_{0},l_{0}]$=[0,3] and its service time to 0. The service time $s_{i}$ for each customer node is set to 0.2. For the time windows for each customer node, we generate them following the methodology outlined in MVMoE (Zhou et al., 2024).
@@
The generation of coordinates for the multiple depots aligns with that of other nodes, with the number of depots set to 3 (Berto et al., 2025). A vehicle is required to return to the same depot it departed from. To guarantee that each starting node is considered in this multi-depot variant, we set the number of trajectories for an instance to be the product of the number of depots and the number of customer nodes, maximizing the exploration of high-quality solutions.
@@
Under the Prize Collecting condition, each node (excluding the depot) is assigned both a prize and a penalty. A prize is collected for each visited node, while a penalty is incurred for each unvisited node. The objective is to minimize the total travel distance plus the sum of penalties for all unvisited nodes. A key constraint is that a minimum amount of total prize must be collected before the vehicle is permitted to return to the depot and terminate its route. In this context, the minimum prize requirement for the four examined problems (PCTSP, SPCTSP, APCTSP, and ASPCTSP) is standardized to 1.0. Consistent with the AM (Kool et al., 2019) framework, the prizes and penalties are generated as follows: $prize\sim\text{Uniform}\left(0,\frac{4}{n}\right)$, and $penalty\sim\text{Uniform}\left(0,3\cdot\frac{k_{n}}{n}\right)$, where $n=100$ and $k_{n}=4$ in our experiments. For the OP, the objective is to maximize collected prizes without exceeding a predefined distance limit (set to 4.0 for OP and 1.0 for AOP), where unvisited nodes carry no penalties. The prize for each node is assigned proportional to its Euclidean distance from the depot. By normalizing distances to the instance’s maximum distance and discretizing them into $[0.01,1.00]$, nodes further from the depot systematically yield higher rewards, with the depot’s prize set to zero.
@@ Appendix C Unified Data Representation
In this section, we provide a detailed description of our UDR $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$ (see Table 9). Each node $\mathbf{u}_{i}$ consists of three components, i.e., $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a positional identifier, unified attribute set, and node-type indicator, respectively. To distinguish linehaul/delivery from backhaul/pickup nodes, we assign negative demand values to backhaul/pickup nodes (positive for linehaul/delivery), making demand the only signed component, following  (Liu et al., 2024; Zhou et al., 2024).
@@ Appendix D Multi-hot Problem Representation
As shown in Table 10, we train URS on eleven classic and varied routing problems to ensure that each feature in the UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$. We adhere to the principle of minimal redundancy and strictly align with existing work on CVRP variant selection (e.g., MVMoE (Zhou et al., 2024) and ReLD (Huang et al., 2025)): most features in UDR appear in at least two training problems, helping prevent URS from overfitting to a single problem. The only exception is the Penalty attribute, as there is structural similarity between the seen PCTSP and unseen SPCTSP. For more detailed analysis and experimental results, please see Appendix J.1.
Then, we evaluate generalization on unseen problems on different combinations of ten constraint attributes (see Appendix B), and the detailed representations of each unseen problem can be found in Table 11 (unseen well-studied VRP variants) and Table 12 (unseen less-explored VRP variants).
@@
ACVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Frequency 2/11 9/11 6/11 2/11 1/11 2/11 2/11 2/11 9/11 2/11 7/11 6/11 2/11
Table 11: The problem representations of unseen $44$ well-studied VRP variants. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
@@
CVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
OCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
MDCVRPBPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
MDOCVRPBPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Table 12: The problem representations of unseen $55$ less-explored VRP variants. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
@@
ACVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
AOCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
AMDCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
AMDOCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
AOP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$
APCTSP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ASPCTSP N/A $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Appendix E Model Architecture and Training
@@ E.1 Encoder
The encoder $\bm{\theta}_{enc}$ with $L$ stacked attention layers transforms $H^{(0)}$ into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. Let ${H}^{(\ell-1)}=\{\mathbf{h}_{i}^{(\ell-1)}\}_{i=0}^{n}$ denote the input to the $\ell$-th attention layer for $\ell=1,\ldots,L$. The outputs for the $i$-th node are computed as follows:
@@
where $\mathrm{IN}(\cdot)$ denotes instance normalization (Ulyanov et al., 2016), $\mathrm{MBM}$ represents the adopted mixed bias module (see Equation 5 for details), and $\mathrm{FF}(\cdot)$ in Equation 12 corresponds to a fully connected neural network with ReLU activation. After $L$ attention layers, the final node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$ encapsulate the advanced feature representations of each node. Notably, to ensure dimensional consistency across bias channels, we normalize $\bm{D}$ and $\bm{D^{\mathrm{T}}}$ by dividing their maximum value before passing the MBM.
@@ E.2 Decoder
$\bm{H}_{\text{dec}}(\bm{\lambda})=((\bm{\lambda}W_{1}+\mathbf{b}_{1})W_{2}+\mathbf{b}_{2})W_{3}+\mathbf{b}_{3},$ (13)
where $W_{1}\in\mathbb{R}^{|\bm{\lambda}|\times d_{h}},\ W_{2}\in\mathbb{R}^{d_{h}\times d_{h}},\ W_{3}\in\mathbb{R}^{d_{h}\times(5\times|\bm{\lambda}|)},\ \mathbf{b}_{1}\in\mathbb{R}^{d_{h}},\ \mathbf{b}_{2}\in\mathbb{R}^{d_{h}},\mathbf{b}_{3}\in\mathbb{R}^{(5\times|\bm{\lambda}|)}$ are learnable parameters, $d_{h}$ is $256$ in this work, following Lin et al. (2022). Next, we use five different linear projections to map the new representation $\bm{h}_{\text{dec}}^{i}(\bm{\lambda})$ of each row to the corresponding decoder parameters. The computation can be expressed as:
@@
where $W^{1}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$, $W^{2}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$, $W^{3}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(1\times d)}$, $W^{4}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$, $W^{5}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$ are learnable parameters. In this way, we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate used linear projection matrices $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ conditioned on the multi-hot problem representation $\bm{\lambda}$, enabling the adaptive construction of high-quality solutions for different problems.